OpenStudy (anonymous):

what # must you add to cmplete the square? x^2+6x=13

OpenStudy (anonymous):

9

OpenStudy (anonymous):

add 22

OpenStudy (anonymous):

i dun think u need to add 22

OpenStudy (anonymous):

it becomes (x+3)^2

OpenStudy (anonymous):

3^2 is just 9

OpenStudy (anonymous):

@techboy to make it a square you should add 22

OpenStudy (anonymous):

reason??

OpenStudy (anonymous):

wat about the 13 on the right hand side have you seen it

OpenStudy (anonymous):

x^2+6x=13 x^2+6x-13=0 (x+3)^2-13-3^2=0 (x+3)^2-22=0 (x+3)^2=22

OpenStudy (anonymous):

u dont add 22 but u get 22 when u add 9